Tickets for a Brittany Spears concert sold for $80 for floor seats and $60 for balcony seats. For one performance, 4,000 tickets were sold, bringing in $288,800. How many of each ticket were sold?
Answer by TimothyLamb(4379)

---
60x + 80y = 288800
x + y = 4000
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = balcony = 1560
y = floor = 2440
---
now by substitution:
---
60x + 80y = 288800
x + y = 4000
---
60x + 80y = 288800
y = 4000 - x
---
60x + 80(4000 - x) = 288800
60x + 80*4000 - 80x = 288800
-20x = -31200
x = -31200/-20
x = 1560
y = 4000 - 1560
y = 2440
---
solution by substitution method:
x = balcony = 1560
y = floor = 2440
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php